OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of 3 and a y-intercept of -2

OpenStudy (anonymous):

when it says the slope is a number like 3 just put a x behind it

OpenStudy (anonymous):

and the y intercept is if it moves up or down so that goes on the end:)

OpenStudy (anonymous):

I'm still really confused haha

OpenStudy (amistre64):

the slope intercept form is called this because it uses the slope and the intercept to form the equation: when given a slope: m and a y intercept: c the slope intercept from is simply: y = mx + c

OpenStudy (anonymous):

mx is the slope right?

OpenStudy (anonymous):

The slope intercept formula is y=mx+b Where only (m) is the slope and (b) is the y intercept For example let's say you have y=6x+4 The slope is (6) and the y intercept is (4). In order to put something in a slope intercept formula you need the slope and the y intercept. In your question you are given the slope 3 and a y-intercept of -2. So all you have to do is plug it in into the formula which is y=mx+b So your answer is y=3x-2. Hope this helps